Suppose that out of the original 250 increase in government spending, 82.5 will be recycled back into purchases of domestically produced goods and services in the second round and 27.225 is spent in the third round. Following this multiplier effect, what value would be recycled in the fourth round of this cycle? Please calculate the final answer to two decimals.
Answer #1

Solution:-

Spending factors = 82.5 / 250 = 0.33 or 33%

SO in fourth round = 27.225 * 0.33

                                     = 8.98425